OpenStudy (anonymous):

solve: (3x-8)(x+1)=(x+1)(x-3)

OpenStudy (anonymous):

oops devide by x+1 both sides so (x is not -1) 3x-8 = x-3 2x = 5 x=2.5

OpenStudy (anonymous):

sorry, i am not solving x I'm solving for 0

OpenStudy (anonymous):

distribute, combine like term, factor 3x^2-8x+3x-8=x^2-3x+1x-3 subtract x^2 on both sides 2x^2-8x+3x-8=-3x+1x-3 2x^2-5x-8=-2x-3 add 2x 2x^2-3x-5 (x+1)(2x-5) set each factor equal to 0